# FSGS: Real-Time Few-shot View Synthesis using Gaussian Splatting

### 01 Dec 2023 v1

@@ 3.4 Optimization
$\mathcal{L}(\boldsymbol{G},\boldsymbol{C})$ $=\lambda_{1}\underbrace{\left\lVert\mathbf{\boldsymbol{C}}-\mathbf{\hat{\boldsymbol{C}}}\right\rVert_{1}}_{\mathcal{L}_{\text{1}}}+\lambda_{2}\underbrace{\operatorname{D-SSIM}({\boldsymbol{C}},\hat{\boldsymbol{C}})}_{\mathcal{L}_{\text{ssim}}}$
$+\lambda_{3}\underbrace{\left\lVert\operatorname{Corr}(\boldsymbol{D}_{\text{ras}},\boldsymbol{D}_{\text{est}})\right\rVert_{1}}_{\mathcal{L}_{\text{prior}}}$ (6)
@@
where $\mathcal{L}_{\text{1}}$, and $\mathcal{L}_{\text{ssim}}$ stands for the photometric loss term between predicted image $\hat{\boldsymbol{C}}$ and ground-truth image $\boldsymbol{C}$. $\mathcal{L}_{\text{prior}}$ represents the geometric regularization term on the training views and synthesized pseudo views. We set $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ as 0.8, 0.2, 0.05 respectively by grid search. The pseudo views sampling is enabled after 2,000 iterations to ensure the Gaussians can roughly represent the scene.
Methods $503\times 381$ Resolution $1006\times 762$ Resolution
@@
RegNeRF 0.21 19.08 0.587 0.336 0.14 18.66 0.535 0.411
FreeNeRF 0.21 19.63 0.612 0.308 0.14 19.13 0.562 0.384
SparseNeRF 0.21 19.86 0.624 0.328 0.14 19.07 0.564 0.392
Ours 458 20.43 0.682 0.248 351 19.71 0.642 0.283
Table 1: Quantitative Comparison in LLFF Datasets, with 3 Training Views. FSGS achieves the best performance in terms of rendering accuracy and inference speed across all resolutions. Significantly, FSGS runs 2,180$\times$ faster than the previous best, SparseNeRF, while improving the SSIM from 0.624 to 0.682, at the resolution of $503\times 381$. We color each cell as best, second best, and third best.
@@
FreeNeRF 0.07 22.78 0.689 0.323 0.04 21.39 0.587 0.377
SparseNeRF 0.07 22.85 0.693 0.315 0.04 21.43 0.604 0.389
Ours 290 23.70 0.745 0.230 203 22.52 0.673 0.313
4 Experiments
4.1 Experimental Settings
LLFF Datasets [29] consist of eight forward-facing real-world scenes. Following RegNeRF [33], we select every eighth image as the test set, and evenly sample sparse views from the remaining images for training purposes. We use 3 views to train all the methods, and evaluate them at resolutions of 1008$\times$756 and 504$\times$378.
Mip-NeRF360 Datasets [2] consist of nine scenes, each featuring a complex central object or area against a detailed background. We utilize the published seven scenes for comparison, using 24 training views with images downsampled to 4$\times$ and 8$\times$. Test images are selected following the same convention with LLFF Datasets. To the best of our knowledge, we are the first to attempt few-shot novel view synthesis in unbounded scenes like Mip-NeRF360. We aim to establish this simple baseline for the challenge of few-shot novel view synthesis in complex large-scale scenarios.
Blender datasets [30] have eight objects with realistic images synthesized by Blender. We align with DietNeRF [23], where we use 8 images for training and 25 for testing. We evalute all methods at resolution of 400$\times$400.
@@ 4.1 Experimental Settings
Mip-NeRF 0.22 20.89 0.830 0.168
3D-GS 332 21.56 0.847 0.130
DietNeRF 0.14 22.50 0.823 0.124
RegNeRF 0.22 23.86 0.852 0.105
FreeNeRF 0.22 24.26 0.883 0.098
SparseNeRF 0.22 24.04 0.876 0.113
Ours 467 24.64 0.895 0.095
Table 3: Quantitative Comparison in Blender Datasets, with 8 Training Views. FSGS outperforms existing few-shot methods and 3D-GS across all metrics, validating the generalization of the proposed techniques to handheld object-level 3D modeling.
@@
Baselines
We compare FSGS with several few-shot NVS methods on these three dataset, including DietNeRF [23], RegNeRF [33], FreeNeRF [55], and SparseNeRF [47]. Additionally, we include comparisons with the high-performing MLP-based NeRF, Mip-NeRF [3], which is primarily designed for dense-view training, and point-based 3D-GS, following its original dense-view training recipe. Following  [33, 50, 26], we report the average PSNR, SSIM, LPIPS scores and FPS for all the methods.
Implementation Details
We implemented FSGS using the PyTorch framework, with initial camera poses and point clouds computed from SfM, based on the specified number of training views. During optimization, we densify the Gaussians every 100 iterations and perform densification after 500 iterations. The total optimization steps are set at 10,000 for all datasets, requiring approximately 9.5 minutes on Blender and LLFF datasets, and $\sim$24 minutes on Mip-NeRF360 datasets. We set proximity threshold $t_{\text{prox}}$ to 10, and the pseudo views are sampled after 2,000 iterations, with $\sigma$ set as 0.1. We utilize the pre-trained Dense Prediction Transformer (DPT) model [37] for zero-shot monocular depth estimation. All results are obtained using an NVIDIA A6000 GPU.
4.2 Comparisons to other Few-shot Methods
Comparisons on LLFF Datasets
As shown in Tab. 1, our method FSGS, despite trained from sparse SfM point clouds, provides the best quantitative results and effectively addresses the insufficient scene coverage in the initialization. Our method surpasses SparseNeRF by 0.57 and 0.64 in PSNR at both test resolutions, while operating 2180 times faster. FSGS also outperforms 3D-GS by 2.60 in PSNR and boost the FPS from 385 to 458, demonstrating that our refined Gaussians are more compact and precise for scene representation from sparse views. The real-time rendering speed makes FSGS a viable choice for practical usages.
The qualitative analysis, as presented in Fig. 5, demonstrates that Mip-NeRF and 3D-GS struggle with the extreme sparse view problem; Mip-NeRF leads to degraded geometric modeling, and 3D-GS tends to produce blurred results in areas of complex geometry. The geometry field regularization in RegNeRF and frequency annealing in FreeNeRF do improve the quality to some extent, but still exhibit insufficient visual quality. SparseNeRF, which also utilizes the monocular depth maps for optimization, manifests as less textural details. In contrast, our proposed Gaussian unpooling, pulls more Gaussians to the unobserved regions and thus recovers more structural details in novel views.
Comparisons on Mip-NeRF360 Datasets
As shown in Tab. 2, methods requiring dense view coverage (Mip-NeRF360, 3D-GS) are outperformed by ours in terms of rendering speed and metrics, across the two resolutions. Methods employing regularizations from their respective geometry and appearance fields (DietNeRF, RegNeRF, FreeNeRF) still suffer from rendering quality. SparseNeRF, by incorporating depth ranking information into NeRF, enhances quality to a degree, but remains far from achieving real-time speed. Our FSGS significantly outperforms NeRF-based approaches, boosting PNSR by 0.85 and improving FPS from 0.07 to 290 at 1/8 resolution. We provide a qualitative comparison in Fig. 9, where we observe that Mip-NeRF360 and SparseNeRF fail to capture the intricate details of scenes and tend to overfit on sparse training views, most notably in areas far away from cameras. In comparison, FSGS recovers the fine-grained details such as the leaves on the ground (Scene: Stump) and the piano keys (Scene: Bonsai), aligning well with the ground truth.
Comparisons on Blender Datasets
@@ 4.2 Comparisons to other Few-shot Methods
Tab. 3 presents the quantitative results on the Blender datasets. Here, our method significantly outperforms the baselines on object-level datasets, with an improvement of 0.40 in PSNR compared to FreeNeRF, although primarily designed for scene-level scenarios with complex geometry. Fig. 7 visualizes the rendered image. We find that DietNeRF hallucinates geometric details, and FreeNeRF exhibits noticeable aliasing effects leading to worse geometry. 3D-GS suffers from excessive blurriness and distorts the edges of the objects. In contrast, our model not only captures the precise geometry of objects but also accurately simulates the shading effects.
@@
Figure 8: Ablation Study by Visualization. 3D-GS [26] (1st row) shows that the baseline method is significantly degraded when the view coverage is insufficient. Gaussian Unpooling provides extra capacity to 3D Gaussians to model the scene, but the learned geometry may not be accurate. Adding Depth Correlation regularization (3rd row) can further improve the modeled details.
@@
17.83 0.582 0.321
18.64 0.603 0.311
19.93 0.644 0.283
20.43 0.682 0.248
Table 4: Ablation Study. Starting from 3D-GS [26] (1st row), we find that our proposed Gaussian Unpooling (2nd row) is more effective than the densification scheme in 3D-GS for few-shot view synthesis. Applying additional supervision from a monocular depth estimator further regularizes the Gaussian optimization towards a better solution (3rd row). Introducing pseudo-view augmentation to apply additional regularization when optimizing Gaussians further enhances the results in a few-shot scenario.
@@ 4.3 Ablation Studies
In Tab. 4, we ablate our design choices on the the LLFF dataset under the 3-view setting.
@@
Effectiveness of Gaussian Unpooling As shown in the second row of Tab. 4, our Gaussian Unpooling expands the scene geometry caused by limited training views, resulting in a PSNR improvement of 0.81 compared to 3D-GS. We also visualize its visual effects in Fig. 8 (2nd row). The heuristic Gaussian densification leads to blurring results, particularly noticeable in areas like bush and grass, our approach enriches structural and visual details.
@@
Impact of Monocular Depth Prior Tab. 4 (3rd row) demonstrates the improvement by introducing depth priors, guiding the Gaussian unpooling towards more plausible geometry. In Fig. 8, we observe that the depth regularization effectively eliminates the artifacts in grassy regions, and enforces more consistent and solid surfaces with geometric coherence. We also display the rendered depth map, where depth regularization leads to depths aligning better with the actual geometric structures.
@@
Pseudo-view Matters in Few-shot Modeling Tab. 4 (4th row) validates the impact of synthesizing more unseen views during training, which anchors the Gaussians to a plausible geometry and further enhances the modeling quality when the geometry in densification is not accurate.
5 Conclusion
@@
In this work, we present a real-time few-shot framework, FSGS, for novel views synthesis within an insufficiently view overlapping. Starting from extremely sparse SfM point cloud, FSGS adopts the point-based representation and proposes an effective Gaussian unpooling method by measuring the proximity of each Gaussian to its neighbor. The adoption of monocular depth priors has proven effective to guide the expanded scene geometry toward a better solution. FSGS is capable of generating photo-realistic images with as few as three images, and perform inference at more than 200FPS, offering new avenues for real-time rendering and more cost-effective capture methods.
A Appendix Part I: More Technical Details
A.1 Initialization
Similar to 3D Gaussian Splatting [26], we start our pipeline from unstructured multi-view images, and calibrate the images using Structure-from-Motion [39]. Next, we will continue the dense stereo matching under COLMAP with the function “patch_match_stereo ” and utilize the fused stereo point cloud from “stereo_fusion ”. We then initialize the SH coefficients at degree 0 and the positions of the 3D Gaussians based on the fused point cloud. Additionally, we remain the rest coefficients and rotation to 0. We also initialize the opacity to 0.1 and set the scale to match the average distance between points.
A.2 Training
During training, we start with a SH degree of 0 for a basic lighting representation, incrementing by 1 every 500 iterations up to a degree of 4 to increase complexity over time. We set the learning rate of position, SH coefficients, opacity, scaling, and rotation to 0.00016, 0.0025, 0.05, 0.005, and 0.001 respectively. At iterations 2000, 5000, and 7000, the opacity for all Gaussians is reset to 0.05 to eliminate the low-opacity floaters. In the Blender dataset [30], the Pearson correlation is only computed in pixels where the depth values are greater than 0. Additionally, we utilize an open-source code [^1] https://pytorch.org/hub/intelisl_midas_v2/ to estimate the inverse depth map for both the input images and the rendered images from pseudo views. We detail the procedures of our proposed FSGS in Algorithm 1.
B Appendix Part II: More Experiments
B.1 Effects of Training Views
We demonstrate the quantitative results of FSGS on LLFF datasets under 3, 6, 9 views in Tab. 5. We can observe that more views consistently exhibit higher visual quality. In Fig.  9, we present a qualitative comparison. We can observe that training with fewer views often leads to overfitting on sparse training data. In contrast, training with 9 views continues to capture the fine-grained structures of the scenes. More views provide a more comprehensive coverage of the scene, capturing more details of the scenes. This richness in data boosts supervision signals during optimization, leading to more detailed and structural texture. Across all the settings, FSGS delivers superior performance compared to all the baselines, affirming the effectiveness of our proposed depth prior.
Figure 9: Qualitative Results on the Effects of Training Views. We visualize the rendered images produced by all methods with training views ranging from three to nine.
Methods PSNR SSIM LPIPS
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
Mip-NeRF 16.10 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.170
3D-GS 17.83 22.87 24.65 0.582 0.732 0.813 0.321 0.204 0.159
DietNeRF 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183
RegNeRF 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161
FreeNeRF 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160
SparseNeRF 19.86 23.64 24.97 0.624 0.784 0.834 0.328 0.202 0.158
Ours 20.31 24.20 25.32 0.682 0.811 0.856 0.248 0.173 0.136
Table 5: Quantitative Analysis on the Effects of Training Views. We conduct experiments by using different training views (from 3 to 9) to test the adopted baseline methods. Our FSGS consistently outperforms other methods across all metrics.
Algorithm 1 The training pipeline of FSGS
1:Training view images =$\mathcal{I}=\{I_{i}\in{}^{H\times W\times 3}\}_{i=1}^{N}$ and their associated camera poses $\mathcal{P}=\{\boldsymbol{\phi}_{i}\in{}^{3\times 4}\}_{i=1}^{N}$.
2:Run SfM with the input images and camera poses and obtain an initial point cloud $\mathcal{P}$, used to define 3D Gaussians function $\mathcal{G}=\{G_{i}(\mu_{i},\sigma_{i},c_{i},\alpha_{i})\}_{i=1}^{K}$.
3:Leverage pretrained depth estimator $\mathcal{E}$ to predict the depth map $D_{i}=\mathcal{E}(I_{i})$.
4:Synthesize pseudo views $\mathcal{{P^{\dagger}}}=\{\boldsymbol{\phi^{\dagger}}_{i}\in{}^{3\times 4}\}_{i=1}^{M}$ from input camera poses $\mathcal{P}$.
5:while until convergence do
6:     Randomly sample an image $I_{i}\in\mathcal{I}$ and the corresponding camera pose $\boldsymbol{\phi}_{i}$
7:     Rasterize the rgb image $\hat{I}_{i}$ and the depth map $\hat{D}_{i}$ with camera pose $\boldsymbol{\phi}_{i}$
8:     $\mathcal{L}=\lambda_{1}\lVert I_{i}-\hat{I}_{i}\rVert_{1}+\lambda_{2}\operatorname{D-SSIM}(I_{i},\hat{I}_{i})+\lambda_{3}\operatorname{Pearson}(D_{i},\hat{D}_{i})$
9:     if $\text{iteration}>t_{iter}$ then
10:         Sample a pseudo camera pose $\boldsymbol{\phi^{\dagger}}_{j}\in\mathcal{P^{\dagger}}$.
11:         Rasterize the rgb image $\hat{I^{\dagger}}_{j}$ and the depth $\hat{D^{\dagger}}_{i}$
12:         Compute the estimated depth as ${D^{\dagger}}_{j}=\mathcal{E}(\hat{I^{\dagger}}_{j})$.
13:         $\mathcal{L}=\mathcal{L}+\lambda_{4}\operatorname{Pearson}({D^{\dagger}}_{j},\hat{D^{\dagger}}_{i})$
14:     end if
15:     if IsRefinement(iteration) then
16:         for $G_{i}(\mu_{i},\sigma_{i},c_{i},\alpha_{i})\in\mathcal{G}$  do
17:              if $\alpha_{i}>\varepsilon$ or IsTooLarge($\mu_{i},\sigma_{i}$then
18:                  RemoveGaussian()
19:              end if
20:              if $\nabla_{p}\mathcal{L}>t_{pos}$ then
21:                  GaussianDensify()
22:              end if
23:              if NoProximity($\mathcal{G}$then
24:                  GaussianUnpooling()
25:              end if
26:         end for
27:     end if
28:     Update Gaussians parameter $\mathcal{G}$ via $\nabla_{\mathcal{G}}\mathcal{L}$.
29:end while
Figure 10: Qualitative Results on Datasets Collected by Mobile Phones. We test the generalization capacity of all methods on self-captured iPhone images, with a calibration process from COLMAP. Our method reveals the majority of scene details despite only adopting three views in training.
Methods 1/8 Resolution 1/4 Resolution
[…]
Mip-NeRF 0.07 14.74 0.337 0.602 0.14 13.84 0.284 0.631
3D-GS 225 16.29 0.408 0.439 127 15.83 0.413 0.530
DietNeRF 0.05 13.62 0.263 0.598 0.08 12.57 0.228 0.722
RegNeRF 0.07 17.41 0.517 0.440 0.14 16.44 0.443 0.504
FreeNeRF 0.07 18.07 0.497 0.426 0.14 17.14 0.462 0.509
SparseNeRF 0.07 18.79 0.539 0.441 0.04 17.82 0.472 0.524
Ours 263 19.54 0.539 0.403 190 18.41 0.493 0.471
Table 6: Quantitative Comparison in Mobile Phone Datasets, with 3 Training Views. Our method continues to achieve the best performance in the challenging mobile phone dataset.
B.2 FSGS on Mobile Phones Data
To validate the generalization capability of FSGS in various real-world settings, we created a new dataset using only a consumer smartphone, the iPhone 15 Pro. This dataset contains three scenes, comprising two indoor scenes and one outdoor scene. Each scene consists of a collection of RGB images under 5712$\times$4284 resolution, with the viewpoint number ranging from 20 to 40. Our data calibration pipeline follows the same process procedures as the LLFF datasets [29], and we also select every 8-th image as the novel views for evaluation. For training, we evenly sample 3 images from the remaining views. These images are then downsampled to 4$\times$ and 8$\times$ for both training and evaluation. We utilize the pretrained monocular depth estimator to predict the depth for input images, and the poses are computed via COLMAP. We compare our method with FreeNeRF [55], SparseNeRF [50], and 3D-GS [26].
Tab.6 presents the quantitative results, where FSGS outperforms SparseNeRF with over 0.75 higher PSNR and runs 3,757× faster, a significant leap that underscores its potential for practical, real-world applications where speed is crucial and the environment is intricate. We also visualize the qualitative results in Fig.10, where FSGS significantly improves the visual quality of the scenes over 3D-GS, particularly in the realm of geometry reconstruction. FreeNeRF and SparseNeRF are constrained by geometric continuity from unseen perspectives and do not fully capitalize on the available depth information.
Figure 11: Visual Comparisons of Predicted Depth. We visualize the estimated scene depth from all baselines. Noticeable NeRF alias artifacts are found in SparseNeRF and FreeNeRF. 3D-GS produces an oversmoothed scene geometry, while our method demonstrates visually pleasing geometric details.
B.3 Visual Comparisons of the Rendered Depth
We demonstrate the qualitative results of the predicted depth for each methods, as shown in Fig. 11. We compare our method with 3DGS [26], FreeNeRF [55] and SparseNeRF [50]. In the left we visualize the ground truth of the images. FSGS significantly outperforms the three baselines in terms of depth quality and details. The depth maps produced by FSGS are not only more accurate but also exhibit a higher level of detail, showcasing its robustness in reconstructing complex scenes. In contrast, both FreeNeRF and SparseNeRF exhibit limitations in geometric modeling and struggle to accurately learn complex geometries, leading to a distorted scene representation. Although SparseNeRF leverage depth prior, it still does not fully capture the fine-grained structures in real-world structures, resulting in a noticeable drop in quality compared to FSGS. 3D-GS, on the other hand, tends to lose fine details in the areas from away the camera, leading to a diminished overall quality in the depth and texture of distant objects.

### 16 Jun 2024 v2

@@ 3.4 Optimization
$\mathcal{L}(\boldsymbol{G},\boldsymbol{C})$ $=\lambda_{1}\underbrace{\left\lVert\mathbf{\boldsymbol{C}}-\mathbf{\hat{\boldsymbol{C}}}\right\rVert_{1}}_{\mathcal{L}_{\text{1}}}+\lambda_{2}\underbrace{\operatorname{D-SSIM}({\boldsymbol{C}},\hat{\boldsymbol{C}})}_{\mathcal{L}_{\text{ssim}}}+\lambda_{3}\underbrace{\left\lVert\operatorname{Corr}(\boldsymbol{D}_{\text{ras}},\boldsymbol{D}_{\text{est}})\right\rVert_{1}}_{\mathcal{L}_{\text{regularization}}}$ (8)
@@
where $\mathcal{L}_{\text{1}}$, and $\mathcal{L}_{\text{ssim}}$ stands for the photometric loss term between predicted image $\hat{\boldsymbol{C}}$ and ground-truth image $\boldsymbol{C}$. $\mathcal{L}_{\text{regularization}}$ represents the geometric regularization term on both the training views and synthesized pseudo views. We set $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ as 0.8, 0.2, 0.05 respectively by grid search. The pseudo views sampling is enabled after 2,000 iterations to ensure the Gaussians can roughly represent the scene.
4 Experiments
4.1 Experimental Settings
LLFF Datasets [29] consist of eight forward-facing real-world scenes. Following RegNeRF [33], we select every eighth image as the test set, and evenly sample sparse views from the remaining images for training. We utilize 3 views to train all the methods, and downsample their resolutions to 4$\times$ and 8$\times$, which are $504\times 378$ and $1008\times 756$ respectively.
Mip-NeRF360 Datasets [3] consist of nine scenes, each featuring a complex central object or area against a detailed background. We utilize 24 training views for comparison, with images downsampled to 4$\times$ and 8$\times$. Test images are selected the same with LLFF Datasets. We aim to establish this challenge benchmark for testing few-shot view synthesis in complex outdoor scenarios.
NeRF-Synthetic Datasets (Blender) [30] have eight objects with realistic images synthesized by Blender. We align with DietNeRF [23], where we use 8 images for training and 25 for testing, at resolution of 400$\times$400.
Shiny Datasets [54] contain more challenging view-dependent effects, like the rainbow reflections on a CD and refraction through a liquid bottle. We evenly select 3 views from the Shiny datasets for training at resolutions of 504×378.
Baselines.
We compare FSGS with several few-shot NVS methods on these three dataset, including DietNeRF [23], RegNeRF [33], FreeNeRF [57], and SparseNeRF [50]. Additionally, we include comparisons with the high-performing Mip-NeRF [2], primarily designed for dense-view training, and point-based 3D-GS, following its original dense-view training recipe. Following  [33, 50, 26], we report the average PSNR, SSIM, LPIPS scores and FPS for all the methods.
Implementation Details.
We implemented FSGS using the PyTorch framework, with initial point cloud computed from SfM, using only the training views. During optimization, we densify the Gaussians every 100 iterations and start densification after 500 iterations. The total optimization steps are set to 10,000, requiring approximately 9.5 minutes on LLFF datasets, and $\sim$24 minutes on Mip-NeRF360 datasets. We set proximity threshold $t_{\text{prox}}$ to 10, and the pseudo views are sampled after 2,000 iterations. We utilize the pre-trained DPT model [37] for depth estimation. All results are obtained using a NVIDIA A6000 GPU.
4.2 Comparisons to other Few-shot Methods
Comparisons on LLFF Datasets.
As shown in Tab. 1, our method FSGS, despite trained from sparse SfM point clouds, provides the best quantitative results and effectively addresses the insufficient scene coverage in the initialization. Our method surpasses SparseNeRF by 0.45dB and 0.81dB in PSNR at both test resolutions, while inferencing 2180 times faster, which makes FSGS a viable choice for practical usages. FSGS also outperforms 3D-GS by 2.88dB in PSNR and boost the FPS from 385 to 458, demonstrating that our refined Gaussians are more compact for scene representation from sparse views.
The qualitative analysis, as presented in Fig. 5, demonstrates that Mip-NeRF and 3D-GS struggle with the extreme sparse view problem; Mip-NeRF [1] leads to degraded geometric modeling, and 3D-GS produces blurred results in areas with complex geometry. The geometry regularization in RegNeRF [32], SparseNeRF [50] and frequency annealing in FreeNeRF [57] do improve the quality to some extent, but still exhibit insufficient visual quality. In contrast, our proposed Proximity-Guided Gaussian Unpooling, and the relative geometric regularizations on both training and synthesized virtual views, pulls more Gaussians to the unobserved regions, and thus recovers more textural and structural details.
Table 1: Quantitative Comparison in LLFF Datasets, with 3 Training Views. FSGS achieves the best performance in terms of rendering accuracy and inference speed across all resolutions. Significantly, FSGS runs 2,180$\times$ faster than the previous best, SparseNeRF, while improving the SSIM from 0.624 to 0.652, at the resolution of $504\times 378$. We color each cell as best, second best, and third best.
Methods 1/8 Resolution 1/4 Resolution
@@ 4.2 Comparisons to other Few-shot Methods
RegNeRF 0.21 19.08 0.587 0.336 0.14 18.06 0.535 0.411
FreeNeRF 0.21 19.63 0.612 0.308 0.14 18.73 0.562 0.384
SparseNeRF 0.21 19.86 0.624 0.328 0.14 19.07 0.564 0.401
Ours 458 20.31 0.652 0.288 351 19.88 0.612 0.340
It is worth to note that, the rendering speed of the 3D Gaussian representation is contingent upon the number of optimized Gaussians, where a reduced Gaussian count leads to faster rendering speed. FSGS outperforms 3D-GS while using less Gaussian counts, by effectively places and optimizes the Gaussians to the most representative positions than 3D-GS. With the same initialization method as 3D-GS, the average optimized Gaussian count from FSGS is 57,513 on LLFF dataset, considerably lower than 63,219 from 3D-GS, which results in a faster rendering speed of FSGS than 3D-GS.
@@
FreeNeRF 0.07 22.78 0.689 0.323 0.04 21.04 0.587 0.377
SparseNeRF 0.07 22.85 0.693 0.315 0.04 21.13 0.600 0.389
Ours 290 23.70 0.745 0.220 203 22.82 0.693 0.293
Comparisons on Mip-NeRF360 Datasets.
As shown in Tab. 2, methods requiring dense view coverage (Mip-NeRF360, 3D-GS) are outperformed by ours in terms of rendering speed and metrics, across the two resolutions. Methods employing regularizations from their respective geometry and appearance fields (DietNeRF, RegNeRF, FreeNeRF, SparseNeRF) still fall short in rendering quality, while remains far from achieving real-time speed. Our FSGS significantly outperforms NeRF-based approaches, boosting PNSR by 0.85dB and improving FPS from 0.07 to 290 at 1/8 resolution. We provide a qualitative comparison in Fig. 6, where we observe that Mip-NeRF360 and SparseNeRF fail to capture the intricate details of scenes and tend to overfit on sparse training views, most notably in areas far away from cameras. In comparison, FSGS recovers the fine-grained details such as the leaves on the ground (Scene: Stump) and the piano keys (Scene: Bonsai), aligning well with the ground truth.
Table 3: Quantitative Comparison in Blender and Shiny Datasets. FSGS outperforms existing few-shot methods and 3D-GS across all metrics, validating the generalization of the proposed techniques to handheld object-level 3D modeling and datasets with challenging reflective effects.
Methods Blender Dataset [30] Shiny Dataset [54]
@@
Mip-NeRF 0.22 20.89 0.830 0.168 0.19 17.37 0.525 0.432
3D-GS 332 21.56 0.847 0.130 316 17.83 0.547 0.385
DietNeRF 0.14 22.50 0.823 0.124 0.16 17.67 0.546 0.403
RegNeRF 0.22 23.86 0.852 0.105 0.19 18.10 0.574 0.378
FreeNeRF 0.22 24.26 0.883 0.098 0.19 18.65 0.586 0.360
SparseNeRF 0.22 24.04 0.876 0.113 0.19 18.81 0.591 0.354
Ours 467 24.64 0.895 0.095 341 19.63 0.612 0.327
@@
Comparisons on Blender Datasets.
@@
The left column of Tab. 3 presents the quantitative results on the Blender datasets. Here, our method significantly outperforms the baselines on object-level datasets, with an improvement of 0.40 in PSNR compared to FreeNeRF, although primarily designed for scene-level scenarios with complex geometry. Fig. 7 visualizes the rendered image. We find that DietNeRF hallucinates geometric details, and FreeNeRF exhibits noticeable aliasing effects. 3D-GS falls short into the excessive blurriness and distorts the edges of the objects. In contrast, our model not only captures the precise geometry of objects but also accurately simulates their shading effects.
Comparisons on Shiny Datasets.
We report the quantitative results of the Shiny datasets on the right column of Tab. 3, which feature complex view-dependent effects. Our method performs better than other baselines and improves the PSNR by 1.80dB compared to 3D-GS. The superior performance validates the robustness of our method to handle unusual and challenging materials, such as CDs or glass.
@@
Figure 8: Ablation Study by Visualization. 3D-GS [26] (1st column) shows that the baseline method is significantly degraded when the view coverage is insufficient. Gaussian Unpooling provides extra capacity to 3D Gaussians to model the scene, but the learned geometry may not be accurate (2nd column). Adding Relative Depth Correspondence regularization (3rd column) can further improve the modeled details.
Table 4: Ablation Study on proposed components. Starting from 3D-GS [26] (1st row), we find that our proposed Gaussian Unpooling (2nd row) is more effective than the densification scheme in 3D-GS for few-shot view synthesis. Applying additional supervision from a monocular depth estimator further regularizes the Gaussian optimization towards a better solution (3rd row). Introducing pseudo-view augmentation to apply additional regularization when optimizing Gaussians further enhances the results in a few-shot scenario.
@@
17.43 0.522 0.321
18.64 0.580 0.311
19.83 0.634 0.297
20.31 0.652 0.288
Table 5: Ablation Study on different depth estimators. We utilize various monocular depth estimators on FSGS, and find that FSGS demonstrates strong robustness over different pretrained depth estimators.
[…]
SparseNeRF 0.21 19.86 0.624 0.328
3D-GS 385 17.43 0.522 0.321
FSGS (MiDaS small) 446 20.17 0.647 0.294
FSGS (DPT Hybrid) 460 20.29 0.652 0.290
FSGS (DPT Large) 458 20.31 0.652 0.288
FSGS (DepthAnything) 468 20.37 0.654 0.289
@@ 4.3 Ablation Studies
We ablate our design choices on the LLFF dataset under the 3-view setting.
Effectiveness of Promity-guided Gaussian Unpooling.
@@
As shown in the second row of Tab. 4, our Promity-guided Gaussian Unpooling expands the scene geometry caused by limited training views, resulting in a PSNR improvement of 1.21dB compared to 3D-GS. We also visualize its visual effects in Fig. 8. The heuristic Gaussian densification leads to blurring results, particularly noticeable in areas like bush and grass, our approach enriches structural and visual details.
Impact of Relative Depth Regularization.
@@
The third row of Tab. 4 demonstrates the improvement by introducing depth priors, guiding the Gaussian unpooling towards more plausible geometry. In Fig. 8, we observe that the depth regularization effectively eliminates the artifacts in grassy regions, and enforces more consistent and solid surfaces with geometric coherence. We also display the rendered depth map, where depth regularization leads to depths aligning better with the actual geometric structures.
Pseudo-view Matters in Few-shot Modeling.
@@
Tab. 4 validates the impact of synthesizing more unseen views during training, which anchors the Gaussians to a plausible geometry and further enhances the modeling quality when the geometry in densification is not accurate.
Robustness on different pretrained depth estimators.
We have shown that FSGS can generalize across various datasets: Blender datasets (object-level), LLFF datasets (indoor), MipNeRF-360 datasets (indoor and outdoor), and self-collected scenes using a mobile phone (Fig.B in the supplementary). To further substantiate its robustness, we employ different monocular depth estimators on LLFF dataset. As shown in Tab. 5, all of the depth estimators outperform the baselines, and the Depth-Anything [58] method achieves the most comparable results. Collectively, FSGS consistently exhibits strong robustness across different pretrained depth estimators.
5 Conclusion and Limitation
@@
In this work, we present a real-time few-shot framework, FSGS, for novel views synthesis within an insufficiently view overlapping. Starting from extremely sparse point clouds, FSGS adopts the point-based representation and proposes an effective Proximity-guided Gaussian Unpooling by measuring the proximity of each Gaussian to its neighbors. The overfitting issue in few-view 3D-GS can be alleviated by the adoption of pseudo-view generation and monocular relative depth correspondences to guide the expanded scene geometry toward a better solution. FSGS is capable of generating photo-realistic images with as few as three images, and perform inference at more than 200FPS, offering new avenues for real-time rendering and more cost-effective capture methods.
Although FSGS notably enhances the quality and efficiency of real-time few-shot neural rendering, it cannot generalize to occluded views that are unobserved during training. We hope that our proposed approach drives new research towards few-shot novel view synthesis in arbitrary 3D scenes.
Acknowledgement
The work is supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.